OpenStudy (anonymous):

Table of Values please help: y= -(x-3)^2+16 i know what to do but i dont think my values are correct

OpenStudy (anonymous):

What values?

OpenStudy (anonymous):

0 25 1 20 2 17 3 16 -1 0 -2 9 -3 20

OpenStudy (anonymous):

OK, u have put in the left hand number (x) and got the right hand number (y). Let's look at the first one, x=0. -(0-3)^2 +16 = -(-3)^2 + 16 = -(-3)(-3) +16 = -9 + 16 = 7. You have calculated 9 + 16 = 25 instead.

OpenStudy (anonymous):

thanks i think I know what my mistake was now

OpenStudy (anonymous):

Easy to do...